# FERMILAB-PUB-24-0776-T

### v1

@@ 1 Introduction
In this manuscript we introduce a new way to test NC NSI in long-baseline accelerator neutrino experiments, combining the effects on neutrino oscillations and scattering rates. Our proposal is to use neutral current event samples in these oscillation experiments to probe NC NSIs, potentially breaking some of the degeneracies present in the other experimental setups. In the presence of NC NSIs with quarks, the neutral current neutrino-nucleus interactions change in a flavor-dependent way, effectively distinguishing the neutrino flavors in the neutral current channel. This flavor non-universality makes oscillation physics relevant in neutral current observables. On top of that, there are nontrivial quantum interference effects involving oscillation amplitudes and flavor-changing scatterings, as we will see later. So far, the NC data set has been used to search for sterile neutrinos at the MINOS and NOvA experiments MINOS:2008ccf ; NOvA:2021smv ; NOvA:2024imi and it has been pointed out that such studies at future experiments can improve over current bounds on sterile neutrinos Gandhi:2017vzo ; Coloma:2017ptb ; Dutta:2019hmb , but to our knowledge there has not been a use of NC data in long-baseline experiments to probe NSIs.
@@ 2 Framework
where $\rho$ is the density matrix describing the neutrino system, and $\hat{\sigma}_{\gamma}$ is a generalized cross section operator which will describe later. The trace is applied to the incoming neutrino flavors, while the different final state neutrino flavor $\gamma$ is summed up incoherently.
@@
where $N_{p},\,N_{n}$ represent the number of free protons and neutrons. It is important to note that the formalism above specifically applies to the simplified scenario of free nucleons at rest, without internal structure. A more precise treatment of NSI effects in the context of coherent elastic neutrino-nucleus scattering is provided in Ref. Coloma:2023ixt . However, the energy spectra in accelerator neutrino experiments typically range from about 0.3 to 10 GeV, a region in which nuclear and nonperturbative effects are relevant (see App. A). To properly account for these, the generalized cross section should be interfaced with the hadronic current in the description of neutrino-nucleus interactions. While a detailed interface within the framework of the Achilles event generator Isaacson:2020wlx ; Isaacson:2022cwh will be left to future work, we propose here the following simplified approach. We use NuWro Juszczak:2005zs ; Golan:2012wx ; Golan:2012rfa to generate standard neutrino neutral current events in the detector material, which for NOvA is CH2. Then we approximate the NC events in the presence of NSIs by
@@
Our approach trivially reduces to the standard cross section when there are no NSIs present, as the prefactor cancels out and $N_{\mathrm{ev}}\propto\sigma_{\mathrm{NuWro}}$. For the case of free nucleons at rest, Eq. (2.13) should, in principle, be exact, although we are simplifying the nucleon as a point particle in this first work. Nevertheless, since we are only considering vector and axial-vector NSIs, we expect this approximation to be reasonably accurate (see Ref. Kopp:2024yvh ). Working within this approximation but at the event level will allow us to better estimate the deposited energy in the neutral current sample.
@@ 4 Results
Having shown the impact of NC NSI on the number of events at long-baseline experiments we now derive constraints on the NSI parameters. We analyze the NOvA data set with 13.6×$\times 10^{20}$ POT from NOvA:2024imi . We assume that the neutrino mass ordering is normal and we use the best fit values of the oscillation parameters from NOvA:2024imi . We perform a ND and FD fit to the data and use a $\chi^{2}$-statistic
@@
Turning now to the results, in Fig. 3, we show the NOvA constraints to non-standard interactions for the vectorial (left) and axial (right) NSI parameters, assuming one NSI parameter is non-zero at a time. For comparison, we also present the respective results from the global fit Coloma:2023ixt . For the vectorial case, the global fit provides marginalized results where all NSI parameters are allowed to vary simultaneously. In contrast, for axial NSIs, the global fit show results obtained by fitting with only one non-zero parameter at a time. We show the global fit results in this case assuming $\varepsilon^{A}_{\alpha\beta}=3\varepsilon^{uA}_{\alpha\beta}=-\varepsilon^{dA}_{\alpha\beta}$, as opposed to what we do here, $\varepsilon^{A}_{\alpha\beta}=3\varepsilon^{uA}_{\alpha\beta}=\varepsilon^{dA}_{\alpha\beta}$. The reason for that is that the global fit has significant contributions from SNO, which is not sensitive to isospin conserving axial NSIs.
@@
Focusing on the vectorial case first, we observe that the global fit constrains the region near zero for all NSI parameters more tightly than current NOvA data. However, there are regions of large NSI values, particularly for $\varepsilon^{V}_{e\mu}$, $\varepsilon^{V}_{\mu\mu}$, and $\varepsilon^{V}_{\tau\tau}$ values near one, which originate from the LMA-Dark region Coloma:2016gei , where the constraints from the global fit are weaker. Our one-parameter-at-a-time fit excludes these regions, potentially contributing valuable information to the global fit. The dominance of the global fit is not unexpected. Since vectorial NSIs change matter effects, therefore modifying oscillation probabilities, there are a large set of experiments that are sensitive to those new physics effects, and there are significant synergies among those experiments. However, only oscillation cannot break the degenerescence of the LMA-Dark solution. This is the role of COHERENT COHERENT:2020iec ; COHERENT:2021xmm and DRESDEN-II Colaresi:2021kus . In particular, since these experiments use neutrinos from pion decay at rest, muon three-body decay or from reactors, they are only sensitive to NSI affecting $\nu_{e}$ and $\nu_{\mu}$. And this is how LBL oscillation experiments are complementary to global fits.
In the axial case, our results are more noteworthy. The NOvA neutral current analysis significantly improves constraints on the NSI parameters $\varepsilon_{\mu\mu}^{A}$, $\varepsilon_{\tau\tau}^{A}$, and $\varepsilon_{e\mu}^{A}$ compared to the global fit, and also helps break degeneracies for $\varepsilon_{e\tau}^{A}$ and $\varepsilon_{\mu\tau}^{A}$. The only parameter where we do not contribute significantly is $\varepsilon_{ee}^{A}$, although we do rule out a region of large NSI values. Here, the global fit is less sensitive to axial NSIs, as these do not affect oscillation probabilities but only NC cross sections. The primary measurement used to probe axial NSIs with quarks comes from a single experiment, SNO SNO:2011hxd , which is sensitive to the axial NC interactions of solar neutrinos with deuterium Bahcall:1988em of the type $\varepsilon^{A}=\varepsilon^{A}_{u}-\varepsilon^{A}_{d}$. Therefore, incorporating neutral current channels in long-baseline experiments would not only improve the sensitivity to general axial NSIs, substantially strenghtening the constraints on these interactions, but also lead to exclusively new bounds for isospin conserving axial NSIs, $\varepsilon^{A}_{u}=\varepsilon^{A}_{d}$.
@@
As mentioned, the constraints presented here are derived by assuming only one non-zero NSI parameter at a time. To appreciate the impact of the degeneracies discussed in the previous section, we show in Fig. 5 the experimental sensitivity to both mock (left) and actual (right) NOvA data in the $\varepsilon^{V}_{\mu\mu}\times\varepsilon^{V}_{\tau\tau}$ plane (upper panels) as well as the $\varepsilon^{A}_{\mu\mu}\times\varepsilon^{A}_{\tau\tau}$ plane (lower panels).
Figure 5: Expected and actual constraints on $\varepsilon_{\mu\mu}^{V}-\varepsilon_{\tau\tau}^{V}$ and $\varepsilon_{\mu\mu}^{A}-\varepsilon_{\tau\tau}^{A}$ assuming the other NSI parameters to be zero and $13.6\times 10^{20}$ POT for NOvA. For the plots on the left we assume a measurement of the SM (left plots), for the plots on the right we analyzed existing NOvA data from Ref. NOvA:2024imi . The best fit values are $\varepsilon^{V}_{\mu\mu}=0.16,~{}\varepsilon^{V}_{\tau\tau}=0.22$ and $\varepsilon^{A}_{\mu\mu}=-0.02,~{}\varepsilon^{A}_{\tau\tau}=0.09$. We took the DUNE sensitivity to vectorial NSIs from Ref. Chatterjee:2021wac .
First, we observe that the degeneracies described in Eqs. (3.3) and (3.4) are clearly visible. For the vectorial case, we also show in the actual data plot the estimated sensitivity of the DUNE experiment to $\varepsilon^{V}_{\mu\mu}$ and $\varepsilon^{V}_{\tau\tau}$ for an exposure of $336$ kton-MW-year Chatterjee:2021wac . As shown, the NOvA neutral current analysis covers a different region of parameter space compared to the future DUNE sensitivity. This highlights the inherent synergy between these two approaches, demonstrating the promising potential of neutral current measurements compared to the more commonly used charged current oscillation analyses in probing NSIs. Furthermore, we anticipate even better sensitivity with the future DUNE neutral current measurements.  [^4] A full analysis of the DUNE neutral current measurement for probing NSIs, accounting for the impact of DUNE-PRISM, is left to future work.
For the axial NSIs, we also observe a degeneracy that is approximately symmetric around the Standard Model expectation, $\varepsilon^{A}_{\mu\mu}=\varepsilon^{A}_{\tau\tau}=0$. This feature is visible in both the mock and actual data analyses.

### v2

@@ 1 Introduction
In this manuscript we introduce a new way to test NC NSI in long-baseline accelerator neutrino experiments, combining the effects on neutrino oscillations and scattering rates. Our proposal is to use neutral current event samples in these oscillation experiments to probe NC NSIs, potentially breaking some of the degeneracies present in the other experimental setups. In the presence of NC NSIs with quarks, the neutral current neutrino-nucleus interactions change in a flavor-dependent way, effectively distinguishing the neutrino flavors in the neutral current channel. This flavor non-universality makes oscillation physics relevant in neutral current observables. On top of that, there are nontrivial quantum interference effects involving oscillation amplitudes and flavor-changing scatterings, as we will see later. So far, the NC data set has been used to search for sterile neutrinos at the MINOS and NOvA experiments MINOS:2008ccf ; NOvA:2021smv ; NOvA:2024imi and it has been pointed out that such studies at future experiments can improve over current bounds on sterile neutrinos Gandhi:2017vzo ; Coloma:2017ptb ; Dutta:2019hmb , but to our knowledge there has not been a use of existing NC data in long-baseline experiments to probe NSIs.
A previous work focused on a potential future DUNE-like experiment Abbaslu:2023vqk and explored NC Deep Inelastic Scattering (DIS) events, deriving bounds on the axial NC NSI of neutrinos with the $u$, $d$, and $s$ quarks. We provide a more comprehensive analysis in this paper. We use existing NOvA data from both near and far detectors (ND and FD) to reduce systematic uncertainties and consider all types of cross sections relevant at the typical NOvA energies.
@@ 2 Framework
where $\rho$ is the density matrix describing the neutrino system, and $\hat{\sigma}_{\gamma}$ is a generalized cross section operator which will be described later. The trace is applied to the incoming neutrino flavors, while the different final state neutrino flavor $\gamma$ is summed up incoherently.
@@
where $N_{p},\,N_{n}$ represent the number of free protons and neutrons. It is important to note that the formalism above specifically applies to the simplified scenario of free nucleons at rest, without internal structure. A more precise treatment of NSI effects in the context of coherent elastic neutrino-nucleus scattering is provided in Ref. Coloma:2023ixt .
However, the energy spectra in accelerator neutrino experiments typically range from about 0.3 to 10 GeV, a region in which nuclear and nonperturbative effects are relevant (see App. A). To properly account for these, the generalized cross section should be interfaced with the hadronic current in the description of neutrino-nucleus interactions. While a detailed interface within the framework of the Achilles event generator Isaacson:2020wlx ; Isaacson:2022cwh will be left to future work, we propose here the following simplified approach. We use NuWro Juszczak:2005zs ; Golan:2012wx ; Golan:2012rfa to generate standard neutrino neutral current events in the detector material, which for NOvA is CH2. Then we approximate the NC events in the presence of NSIs by
@@
Our approach trivially reduces to the standard cross section when there are no NSIs present, as the prefactor cancels out and $N_{\mathrm{ev}}\propto\sigma_{\mathrm{NuWro}}$. For the case of free nucleons at rest, Eq. (2.13) should, in principle, be exact, although we are simplifying the nucleon as a point particle in this first work. Nevertheless, since we are only considering vector and axial-vector NSIs, we expect this approximation to be reasonably accurate (see Ref. Kopp:2024yvh ). However, note that deviations from our approach are expected when taking form factors into account for example due to the specific flavor dependence of various inelastic processes, from the phase space of a final state with multiple particles, and the effect of strange quarks in the nucleon Ilma:2024lkp .
In Ilma:2024lkp it has been shown that in the presence of axial NSI the elastic NC cross section depends on the isoscalar form factor of the nucleon which is not probed in SM processes, demonstrating that the inclusion of form factors in this analysis will further probe our understanding of nuclear effects, beyond the present study.
Working within this approximation but at the event level will allow us to better estimate the deposited energy in the neutral current sample.
@@ 4 Results
Having shown the impact of NC NSI on the number of events at long-baseline experiments we now derive constraints on the NSI parameters. We analyze the NOvA data set with 13.6×$\times 10^{20}$ POT from NOvA:2024imi . We assume that the neutrino mass ordering is normal and we use the best fit values of the oscillation parameters from NOvA:2024imi which we keep fixed. A different choice for the true values of oscillation parameters will not affect our results much. We perform a ND and FD fit to the data and use a $\chi^{2}$-statistic
@@
Turning now to the results, in Fig. 3, we show the NOvA constraints to non-standard interactions for the vectorial (left) and axial (right) NSI parameters, assuming one NSI parameter is non-zero at a time. For comparison, we also present the respective results from the global fit Coloma:2023ixt . To allow a comparison with the global fit results we rescale the global fit results bounds as $\epsilon_{\alpha\beta}^{(A,V)}=3\epsilon_{\alpha\beta}^{uA(uV)}$. For the vectorial case, the global fit provides marginalized results where all NSI parameters are allowed to vary simultaneously. Although this hinders a direct comparison with our results and the global fit, our choice of considering only one non-zero NSI parameter-at-a-time is motivated due to several degeneracies in the NSI parameter space which make it challenging for one experiment to derive constraints on one NSI parameter allowing all of them to be non-zero. Note that global fits overcome this challenge by considering several experiments simultaneously to break these degeneracies. The inclusion of NC NSI detection data, as proposed in this manuscript, can provide a valuable opportunity to break remaining degeneracies in the global fit. However a quantitative estimate of this is beyond the scope of this work. In contrast, for axial NSIs, the global fit show results obtained by fitting with only one non-zero parameter at a time.
@@
Focusing on the vectorial case first, we observe that the global fit constrains the region near zero for all NSI parameters more tightly than current NOvA data. However, there are regions of large NSI values, particularly for $\varepsilon^{V}_{e\mu}$, $\varepsilon^{V}_{\mu\mu}$, and $\varepsilon^{V}_{\tau\tau}$ values near one, which originate from the LMA-Dark region Coloma:2016gei , where the constraints from the global fit are weaker. Our one-parameter-at-a-time fit excludes these regions, potentially contributing valuable information to the global fit. The dominance of the global fit is not unexpected. Since vectorial NSIs change matter effects, therefore modifying oscillation probabilities, there are a large set of experiments that are sensitive to those new physics effects, and there are significant synergies among those experiments. However, only oscillation cannot break the degeneracy of the LMA-Dark solution. This is the role of COHERENT COHERENT:2020iec ; COHERENT:2021xmm and DRESDEN-II Colaresi:2021kus . In particular, since these experiments use neutrinos from pion decay at rest, muon three-body decay or from reactors, they are only sensitive to NSI affecting $\nu_{e}$ and $\nu_{\mu}$. And this is how LBL oscillation experiments are complementary to global fits.
In the axial case, our results are more noteworthy. The NOvA neutral current analysis significantly improves constraints on the NSI parameters $\varepsilon_{\mu\mu}^{A}$, $\varepsilon_{\tau\tau}^{A}$, and $\varepsilon_{e\mu}^{A}$ compared to the global fit, and also helps break degeneracies for $\varepsilon_{e\tau}^{A}$ and $\varepsilon_{\mu\tau}^{A}$. The only parameter where we do not contribute significantly is $\varepsilon_{ee}^{A}$, although we do rule out a region of large NSI values. Here, the global fit is less sensitive to axial NSIs, as these do not affect oscillation probabilities but only NC cross sections (see also Coloma:2024ict ). The primary measurement used to probe axial NSIs with quarks comes from a single experiment, SNO SNO:2011hxd , which is sensitive to the axial NC interactions of solar neutrinos with deuterium Bahcall:1988em of the type $\varepsilon^{A}=\varepsilon^{A}_{u}-\varepsilon^{A}_{d}$. Therefore, incorporating neutral current channels in long-baseline experiments would not only improve the sensitivity to general axial NSIs, substantially strenghtening the constraints on these interactions, but also lead to exclusively new bounds for isospin conserving axial NSIs, $\varepsilon^{A}_{u}=\varepsilon^{A}_{d}$.
@@
As mentioned, the constraints presented here are derived by assuming only one non-zero NSI parameter at a time. To appreciate the impact of the degeneracies discussed in the previous section, we show in Fig. 5 the experimental sensitivity to both mock (left) and actual (right) NOvA data in the $\varepsilon^{V}_{\mu\mu}\times\varepsilon^{V}_{\tau\tau}$ plane (upper panels) as well as the $\varepsilon^{A}_{\mu\mu}\times\varepsilon^{A}_{\tau\tau}$ plane (lower panels). We have kept the true values of the oscillation parameters fixed to the values from NOvA:2024imi .
Figure 5: Expected and actual constraints on $\varepsilon_{\mu\mu}^{V}-\varepsilon_{\tau\tau}^{V}$ and $\varepsilon_{\mu\mu}^{A}-\varepsilon_{\tau\tau}^{A}$ assuming the other NSI parameters to be zero and $13.6\times 10^{20}$ POT for NOvA. For the plots on the left we assume a measurement of the SM (left plots), for the plots on the right we analyzed existing NOvA data from Ref. NOvA:2024imi . For all plots we fix the true values of the oscillation parameters on the values from NOvA:2024imi . The best fit values are $\varepsilon^{V}_{\mu\mu}=0.16,~{}\varepsilon^{V}_{\tau\tau}=0.22$ and $\varepsilon^{A}_{\mu\mu}=-0.02,~{}\varepsilon^{A}_{\tau\tau}=0.09$. We took the DUNE sensitivity to vectorial NSIs from Ref. Chatterjee:2021wac .
First, we observe that the degeneracies described in Eqs. (3.3) and (3.4) are clearly visible. For the vectorial case, we also show in the actual data plot the estimated sensitivity of the DUNE experiment to $\varepsilon^{V}_{\mu\mu}$ and $\varepsilon^{V}_{\tau\tau}$ for an exposure of $336$ kton-MW-year Chatterjee:2021wac . As shown, the NOvA neutral current analysis covers a different region of parameter space compared to the future DUNE sensitivity. This highlights the inherent synergy between these two approaches, demonstrating the promising potential of neutral current measurements compared to the more commonly used charged current oscillation analyses in probing NSIs. Furthermore, we anticipate even better sensitivity with the future DUNE neutral current measurements Abbaslu:2023vqk  [^4] A full analysis of the DUNE neutral current measurement for probing NSIs, accounting for the impact of DUNE-PRISM, is left to future work. For axial NSIs, we also observe a degeneracy that is approximately symmetric around the Standard Model expectation, $\varepsilon^{A}_{\mu\mu}=\varepsilon^{A}_{\tau\tau}=0$. This feature is visible in both the mock and actual data analyses.
@@
Finally, we mention that if NC NSI is realized in nature, adding the NC detection channel to global fits when extracting the oscillation parameters will be an interesting possibility, as NC events are considered background to the CC channel, and can add valuable information as well as aid in breaking degeneracies between multiple NSI parameters.